On a given day Adam enjoys the consumption of beer (B) and wine (W) according to the function:

U(B, W) = 20B - B2 +18W -3W2

How many glasses of beer and glasses of wine does he consume during a day? How much utility does he get from this consumption? (Cost is no object to Adam)
Answer #1

Maximize the utility by keeping the partial derivatives equal to zero

U'B = 20 - 2B = 0 which gives B = 10

U'W = 18 - 6W = 0 which gives W = 3

Therefore we should consume 10 glasses of beer and 3 glasses of wine

Utility = 20*10-10^2+18*3-3*3^2

= 127 utils